OpenStudy (anonymous):

ln e^x+4y ??? im confused

ganeshie8 (ganeshie8):

why ?

OpenStudy (anonymous):

because i forgot how to do it

ganeshie8 (ganeshie8):

what exactly are you trying to do ?

OpenStudy (anonymous):

simplify

ganeshie8 (ganeshie8):

Ohkay \[\large \ln e^{x+4y}\] like this ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

use below log property : \[\ln (a^b) = b\times \ln(a)\]

ganeshie8 (ganeshie8):

\[\large \ln (e^{x+4y}) = (x+4y) \times \ln (e)\]

OpenStudy (anonymous):

what do i do ? idk how to begin

OpenStudy (anonymous):

ln(e) = 1

OpenStudy (anonymous):

so it means that ln cancels e

OpenStudy (anonymous):

so is just the x+4y?

OpenStudy (anonymous):

so you end up with just x+4y

ganeshie8 (ganeshie8):

\[\large \ln (e^{x+4y}) = (x+4y) \times \ln (e) = (x+4y)\times 1 = x+4y\] yes thats the end of simplification. There is a nice way to make sense of these identties by using the fact that \(\ln (x)\) and \(e^x\) are inverses of each other. Are you good with inverse functions and graphs ?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

\[\ln(e ^{x+4y})=(x+4y) \]

OpenStudy (anonymous):

yeah what @ganeshie8 says graph ln( and e^x so you can see it

ganeshie8 (ganeshie8):

its easy :